OpenStudy (anonymous):

Please help me to verify tan(3x)-tan(2x)-tan(x)=tan(3x)tan(2x)tan(x) is true.And this is an identity.

OpenStudy (loser66):

\[tan(3x-2x) = \frac{tan 3x-tan2x}{1+tan3xtan2x}\\tan3x-tan2x = tanx(1+tan3xtan2x)=tanx + tan3xtan2xtanx\] then the whole thing minus tan x to get tan (3x)tan(2x) tan x. that's it

OpenStudy (anonymous):

Thank you i appreciate your help.

OpenStudy (loser66):

yw